# RapVerse: Coherent Vocals and Whole-Body Motions Generations from Text

### v1

@@ Abstract
In this work, we introduce a challenging task for simultaneously generating 3D holistic body motions and singing vocals directly from textual lyrics inputs, advancing beyond existing works that typically address these two modalities in isolation. To facilitate this, we first collect the RapVerse dataset, a large dataset containing synchronous rapping vocals, lyrics, and high-quality 3D holistic body meshes. With the RapVerse dataset, we investigate the extent to which scaling autoregressive multimodal transformers across language, audio, and motion can enhance the coherent and realistic generation of vocals and whole-body human motions. For modality unification, a vector-quantized variational autoencoder is employed to encode whole-body motion sequences into discrete motion tokens, while a vocal-to-unit model is leveraged to obtain quantized audio tokens preserving content, prosodic information and singer identity. By jointly performing transformer modeling on these three modalities in a unified way, our framework ensures a seamless and realistic blend of vocals and human motions. Extensive experiments demonstrate that our unified generation framework not only produces coherent and realistic singing vocals alongside human motions directly from textual inputs, but also rivals the performance of specialized single-modality generation systems, establishing new benchmarks for joint vocal-motion generation. The project page is available for research purposes at https://vis-www.cs.umass.edu/RapVerse.
@@ 1 Introduction
In the evolving landscape of multi-modal content generation in terms of sound and motion, significant strides have been made in individual modalities, including text-to-music schneider2023mo ; agostinelli2023musiclm ; huang2023noise2music , text-to-vocal liu2022diffsinger , text-to-motion guo2022generating ; zhang2023t2m ; chen2023executing ; jiang2024motiongpt ; lu2023humantomato , and audio-to-motion yi2023generating ; habibie2021learning ; liu2023emage generation. These developments have paved the way for creating more dynamic and interactive digital content. Despite these advancements, existing works predominantly operate in silos, addressing each modality in isolation. However, there’s strong psychological evidence that for human beings, the generation of sound and motion are highly related and coupled lakoff1980metaphors . A unified system for joint generation allows for a more expressive and nuanced communication of emotions, intentions, and context, where the generation of one modality could guide and assist the other in a coherent and efficient way.
@@
An important question naturally arises: what constitutes a good model for unified generation of sound and motion? Firstly, we consider textual lyrics as the proper form of inputs for the unified system, since text provides a highly expressive, interpretable flexible means of conveying information by human beings, and could serve as a bridge between various modalities. Previous efforts explore scores liu2022diffsinger , action commands zhang2023t2m ; chen2023executing ; jiang2024motiongpt , or audio signals yi2023generating as inputs, which are inferior to textual inputs in terms of semantic richness, expressiveness and flexible integration of different modalities.
@@
To build such a joint generation system, the primary challenges include: 1) the scarcity of datasets that provide lyrics, vocals, and 3D whole-body motion annotations simultaneously; and 2) the need for a unified architecture capable of coherently synthesizing vocals and motions from text. In response to these challenges, we have curated RapVerse, a large-scale dataset featuring a comprehensive collection of lyrics, singing vocals, and 3D whole-body motions. Despite the existence of datasets available for text-to-vocal liu2022diffsinger ; huang2021multi ; duan2013nus ; sharma2021nhss , text-to-motion plappert2016kit ; mahmood2019amass ; guo2022generating ; lin2024motion , and audio-to-motion cao2017realtime ; habibie2021learning ; ginosar2019learning ; fanelli20103 ; cudeiro2019capture ; wuu2022multiface , the landscape lacks a unified dataset that encapsulates singing vocals, whole-body motion, and lyrics simultaneously. Most notably, large text-to-vocal datasets huang2021multi ; zhang2022m4singer are predominantly in Chinese, limiting their applicability for English language research and lacking any motion data. And text-to-motion datasets plappert2016kit ; guo2022generating ; lin2024motion typically focus on text descriptions of specific actions paired with corresponding motions without audio data, often not covering whole body movements. Moreover, audio-to-motion datasets liu2022diffsinger ; lu2023co focus primarily on speech rather than singing. A comparison of existing related datasets is demonstrated in Table. 1. The RapVerse dataset is divided into two distinctive parts to cater to a broad range of research needs: 1) a Rap-Vocal subset containing a large number of pairs of vocals and lyrics, and 2) a Rap-Motion subset encompassing vocals, lyrics, and human motions. The Rap-Vocal subset contains 108.44 hours of high-quality English singing voice in the rap genre without background music. Paired lyrics and vocals are crawled from the Internet from 32 singers, with careful cleaning and post-processing. On the other hand, the Rap-Motion subset contains 26.8 hours of rap performance videos with 3D holistic body mesh annotations in SMPL-X parameters pavlakos2019expressive using the annotation pipeline of Motion-X lin2024motion , synchronous singing vocals and corresponding lyrics.
@@
With the RapVerse dataset, we explore how far we can push by simply scaling autoregressive multimodal transformers with language, audio, and motion for a coherent and realistic generation of vocals and whole-body human motions. To this end, we unify different modalities as token representations. Specifically, three VQVAEs van2017neural are utilized to compress whole-body motion sequences into three-level discrete tokens for head, body, and hand, respectively. For vocal generation, previous works min2021meta ; donahue2020end ; liu2022diffsinger ; min2021meta share a common paradigm, producing mel-spectrograms of audio signals from input textual features and additional music score information, following with a vocoder oord2016wavenet ; valin2019lpcnet ; yang2021multi to reconstruct the phase. We draw inspiration from the speech resynthesis domain polyak2021speech , and learn a self-supervised discrete representation to quantize raw audio signal into discrete tokens while preserving the vocal content and prosodic information. Then, with all the inputs in discrete representations, we leverage a transformer to predict the discrete codes of audio and motion in an autoregressive fashion. Extensive experiments demonstrate that this straightforward unified generation framework not only produces realistic singing vocals alongside human motions directly from textual inputs but also rivals the performance of specialized single-modality generation systems.
@@
  • We release RapVerse, a large dataset featuring synchronous singing vocals, lyrics, and high-quality 3D holistic SMPL-X parameters.
@@ 2.1 Text to Vocal Generation
Text-to-audio Dataset. There exists several singing vocal datasets, yet they each has constraints. For instance, PopCS liu2022diffsinger and OpenSinger huang2021multi are limited to Chinese, while NUS-48E duan2013nus and NHSS sharma2021nhss have only a few hours of songs. Nevertheless, JVS-MuSiC tamaru2020jvs and NUS-48E duan2013nus offer a few hours of songs from dozens of singers, whereas OpenSinger huang2021multi provides a more extensive collection with tens of hours from a single singer. Notably, our dataset represents the first to be specifically curated for rap songs of multiple singers with 108 hours.
@@
Text-to-vocal Models. Recent advancements in text-to-speech (TTS) models, including WavNet oord2016wavenet , FastSpeech 1 and 2 ren2019fastspeech ; ren2020fastspeech , and EATS donahue2020end , have significantly improved the quality of synthesized speech. However, singing voice synthesis (SVS) presents a greater challenge due to its reliance on additional musical scores and lyrics. Recent generation models liu2022diffsinger ; he2023rmssinger ; he2023rmssinger ; zhang2022wesinger ; kim2023muse perform excellently in generating singing voices. However, their performance tends to deteriorate when encountering out-of-distribution data. To handle this, StyleSinger zhang2023stylesinger introduces a Residual Style Adaptor and an Uncertainty Modeling Layer Normalization to handle this.
2.2 Text to Motion Generation
Text-to-motion Dataset. Current text-motion datasets, such as KIT plappert2016kit , AMASS mahmood2019amass , and HumanML3D guo2022generating , are constrained by limited data coverage, typically spanning only tens of hours and lacking whole-body motion representation. To overcome these shortcomings, Motion-X lin2024motion was developed, providing an extensive dataset that encompasses whole-body motions.
Text-to-motion Models. Text-to-motion models petrovich2022temos ; zhang2024motiondiffuse ; guo2022tm2t ; guo2022generating ; lu2023humantomato ; chen2023executing have gained popularity for their user convenience. Recent advancements have focused on diffusion models zhang2024motiondiffuse ; kim2023flame ; tevet2022human , which, unlike deterministic generation models ahuja2019language2pose ; ghosh2021synthesis , enable finer-grained and diverse output generation. Chen et al. chen2023executing proposed a motion latent-based diffusion model to enhance generative quality and reduce computational costs. However, the misalignment between natural language and human motions presents challenges. Lu et al. lu2023humantomato introduced the Text-aligned Whole-body Motion generation framework to address these issues by generating high-quality, diverse, and coherent facial expressions, hand gestures, and body motions simultaneously.
2.3 Audio to Motion Generation
Audio-to-motion Dataset. Co-Speech Datasets can be classified into 1: pseudo-labeled (PGT) and 2: motion-captured (mocap). PGT datasets cao2017realtime ; habibie2021learning ; ginosar2019learning ; yi2023generating utilize disconnected 2D or 3D key points to represent the body, allowing extracting at a lower cost but with limited accuracy. On the other hand, mocap datasets provide annotations for limited body parts, with some focusing solely on the head fanelli20103 ; cudeiro2019capture ; wuu2022multiface or the body takeuchi2017creating ; ferstl2018investigating . Differing from them, BEATX liu2023emage contains mesh data of both head and body. However, these datasets typically focus on speech to motion generation. Our RapVerse, in contrast, contains paired text-vocal-motion data, enabling simultaneous motion and vocal generation.
Audio-to-motion Models. Audio-to-motion models taylor2017deep ; shen2023difftalk ; zhang2023sadtalker ; liu2023emage ; yi2023generating aim to produce human motion from audio innputs. Recognizing the intricate relationship between audio and human face, TalkSHOW yi2023generating separately generates face parts and other body parts. However, this approach exhibits limitations, such as the absence of generating lower body and global motion. EMAGE liu2023emage utilizes masked gestures together with vocals for a generation. Our model, however, goes a step further by being the first to generate paired audio-motion data directly from text.
Dataset Language ##\##Singers Hours
NUS-48E duan2013nus English 12 1.91
NHSS sharma2021nhss English 10 7
JVS-MuSiC tamaru2020jvs Japanese 100 2.28
Tohoku Kiritan ogawa2021tohoku Japanese 1 1
PopCS liu2022diffsinger Chinese 1 5.89
OpenSinger huang2021multi Chinese 66 50
Opencpop wang2022opencpop Chinese 1 5.25
M4Singer zhang2022m4singer Chinese 20 29.77
@@ 2.3 Audio to Motion Generation
AMASS mahmood2019amass Text
@@
HumanML3D guo2022generating Text
@@
Motion-X lin2024motion Text
@@
Speech2Gesture ginosar2019learning Audio
@@
Talkshow yi2023generating Audio
@@
(b) Motion Datasets
Table 1: Comparison of Audio and Motion Datasets. (a) compares our Rap-Vocal Subset with existing singing vocal datasets. (b) compares our Rap-Motion Subset with existing human motion datasets. : “Whole” means whole-body motion including face, body, and hand.
@@ 3.1 Rap-Vocal Subset
Vocal and Background Music Separation. Since the crawled songs are mixed with rapping vocals and background music, and we aim to synthesize singing vocals from separated clean data, we utilize Spleeter hennequin2020spleeter , the state-of-the-art open-source vocal-music source separation tool to separate and extract rap vocal voices and accompanying background music from the collected songs. Following ren2020deepsinger , we normalize the loudness of the vocal voices to a fixed loudness level.
@@ 3.2 Rap-Motion Subset
The Rap-Motion subset contains 26.8 hours of rap performance videos with 3D holistic body mesh annotations in SMPL-X parameters pavlakos2019expressive , synchronous singing vocals, and corresponding lyrics. We introduce the collection pipeline of this subset as follows.
@@
Audio Data Processing. Similar to the Rap-Vocal subset, we leverage Spleeter spleeter2020 to isolate singing vocals from accompanying music. Given that YouTube videos typically lack paired lyrics, we employ an ASR model, Whisper radford2023robust , to accurately transcribe vocals into corresponding text.
Video Data Processing. To ensure the collection of high-quality video clips for motion annotation, we implemented a semi-automatic process to filter out undesirable content, such as advertisements, transition frames, changes in shots, and flashing lights. Initially, we applied YOLO redmon2018yolov3 for human detection to discard frames where no humans were detected. Subsequently, we utilized RAFT teed2020raft to assess the motion magnitude, employing a threshold to eliminate frames affected by camera instability. We then perform meticulous manual curation on the extracted clips, retaining only those of the highest quality. Finally, we follow the pipeline of the optimized-based method Motion-X lin2024motion to extract 3D whole-body meshes from monocular videos. Specifically, we adopt the SMPL-X pavlakos2019expressive for motion representations, given a T-frame video clip, the corresponding pose states $\mathcal{M}$ are represented as:
=$\mathcal{M}=\{\mathcal{M}_{f},\mathcal{M}_{b},\mathcal{M}_{h},\zeta,\epsilon\}$ (1)

### v3

# RapVerse: Coherent Vocals and Whole-Body Motion Generation from Text

@@ Abstract
In this work, we introduce a challenging task for simultaneously generating 3D holistic body motions and singing vocals directly from textual lyrics inputs, advancing beyond existing works that typically address these two modalities in isolation. To facilitate this, we first collect the RapVerse dataset, a large dataset containing synchronous rapping vocals, lyrics, and high-quality 3D holistic body meshes. With the RapVerse dataset, we investigate the extent to which scaling autoregressive multimodal transformers across language, audio, and motion can enhance the coherent and realistic generation of vocals and whole-body human motions. For modality unification, a vector-quantized variational autoencoder is employed to encode whole-body motion sequences into discrete motion tokens, while a vocal-to-unit model is leveraged to obtain quantized audio tokens preserving content, prosodic information and singer identity. By jointly performing transformer modeling on these three modalities in a unified way, our framework ensures a seamless and realistic blend of vocals and human motions. Extensive experiments demonstrate that our unified generation framework not only produces coherent and realistic singing vocals alongside human motions directly from textual inputs, but also rivals the performance of specialized single-modality generation systems, establishing new benchmarks for joint vocal-motion generation. Video demonstration and more information can be found on the project page [^1] https://jiabenchen.github.io/RapVerse/..
@@ 1 Introduction
In the evolving landscape of multi-modal content generation in terms of sound and motion, significant strides have been made in individual modalities, including text-to-music [schneider2023mo, agostinelli2023musiclm, huang2023noise2music], text-to-vocal [liu2022diffsinger], text-to-motion [guo2022generating, zhang2023t2m, chen2023executing, jiang2024motiongpt, lu2023humantomato], and audio-to-motion [yi2023generating, habibie2021learning, liu2023emage] generation. These developments have paved the way for creating more dynamic and interactive digital content. Despite these advancements, existing works predominantly operate in silos, addressing each modality in isolation. However, there’s strong psychological evidence that for human beings, the generation of sound and motion are highly related and coupled [lakoff1980metaphors]. A unified system for joint generation allows for a more expressive and nuanced communication of emotions, intentions, and context, where the generation of one modality could guide and assist the other in a coherent and efficient way.
@@
An important question naturally arises: what constitutes a good model for unified generation of sound and motion? Firstly, we consider textual lyrics as the proper form of inputs for the unified system, since text provides a highly expressive, interpretable flexible means of conveying information by human beings, and could serve as a bridge between various modalities. Previous efforts explore scores [liu2022diffsinger], action commands [zhang2023t2m, chen2023executing, jiang2024motiongpt], or audio signals [yi2023generating] as inputs, which are inferior to textual inputs in terms of semantic richness, expressiveness and flexible integration of different modalities.
@@
To build such a joint generation system, the primary challenges include: 1) the scarcity of datasets that provide lyrics, vocals, and 3D whole-body motion annotations simultaneously; and 2) the need for a unified architecture capable of coherently synthesizing vocals and motions from text. In response to these challenges, we have curated RapVerse, a large-scale dataset featuring a comprehensive collection of lyrics, singing vocals, and 3D whole-body motions. Despite the existence of datasets available for text-to-vocal [liu2022diffsinger, huang2021multi, duan2013nus, sharma2021nhss], text-to-motion [plappert2016kit, mahmood2019amass, guo2022generating, lin2024motion], and audio-to-motion [cao2017realtime, habibie2021learning, ginosar2019learning, fanelli20103, cudeiro2019capture, wuu2022multiface], the landscape lacks a unified dataset that encapsulates singing vocals, whole-body motion, and lyrics simultaneously. Most notably, large text-to-vocal datasets [huang2021multi, zhang2022m4singer] are predominantly in Chinese, limiting their applicability for English language research and lacking any motion data. And text-to-motion datasets [plappert2016kit, guo2022generating, lin2024motion] typically focus on text descriptions of specific actions paired with corresponding motions without audio data, often not covering whole body movements. Moreover, audio-to-motion datasets [liu2022diffsinger, lu2023co] focus primarily on speech rather than singing. A comparison of existing related datasets is demonstrated in Table. 1. The RapVerse dataset is divided into two distinctive parts to cater to a broad range of research needs: 1) a Rap-Vocal subset containing a large number of pairs of vocals and lyrics, and 2) a Rap-Motion subset encompassing vocals, lyrics, and human motions. The Rap-Vocal subset contains 108.44 hours of high-quality English singing voice in the rap genre without background music. Paired lyrics and vocals are crawled from the Internet from 32 singers, with careful cleaning and post-processing. On the other hand, the Rap-Motion subset contains 26.8 hours of rap performance videos with 3D holistic body mesh annotations in SMPL-X parameters [pavlakos2019expressive] using the annotation pipeline of Motion-X [lin2024motion], synchronous singing vocals and corresponding lyrics.
@@
With the RapVerse dataset, we explore how far we can push by simply scaling autoregressive multimodal transformers with language, audio, and motion for a coherent and realistic generation of vocals and whole-body human motions. To this end, we unify different modalities as token representations. Specifically, three VQVAEs [van2017neural] are utilized to compress whole-body motion sequences into three-level discrete tokens for head, body, and hand, respectively. For vocal generation, previous works [min2021meta, donahue2020end, liu2022diffsinger, min2021meta] share a common paradigm, producing mel-spectrograms of audio signals from input textual features and additional music score information, following with a vocoder [oord2016wavenet, valin2019lpcnet, yang2021multi] to reconstruct the phase. We draw inspiration from the speech resynthesis domain [polyak2021speech], and learn a self-supervised discrete representation to quantize raw audio signal into discrete tokens while preserving the vocal content and prosodic information. Then, with all the inputs in discrete representations, we leverage a transformer to predict the discrete codes of audio and motion in an autoregressive fashion. Extensive experiments demonstrate that this straightforward unified generation framework not only produces realistic singing vocals alongside human motions directly from textual inputs but also rivals the performance of specialized single-modality generation systems.
@@
  • We collect RapVerse, a large dataset featuring synchronous singing vocals, lyrics, and high-quality 3D holistic SMPL-X parameters.
@@ 2.1 Text to Vocal Generation
Text-to-audio Dataset. There exists several singing vocal datasets, yet they each has constraints. For instance, PopCS [liu2022diffsinger] and OpenSinger [huang2021multi] are limited to Chinese, while NUS-48E [duan2013nus] and NHSS [sharma2021nhss] have only a few hours of songs. Nevertheless, JVS-MuSiC [tamaru2020jvs] and NUS-48E [duan2013nus] offer a few hours of songs from dozens of singers, whereas OpenSinger [huang2021multi] provides a more extensive collection with tens of hours from a single singer. Notably, our dataset represents the first to be specifically curated for rap songs of multiple singers with 108 hours.
@@
Text-to-vocal Models. Recent advancements in text-to-speech (TTS) models, including WavNet [oord2016wavenet], FastSpeech 1 and 2 [ren2019fastspeech, ren2020fastspeech], and EATS [donahue2020end], have significantly improved the quality of synthesized speech. However, singing voice synthesis (SVS) presents a greater challenge due to its reliance on additional musical scores and lyrics. Recent generation models [liu2022diffsinger, he2023rmssinger, he2023rmssinger, zhang2022wesinger, kim2023muse] perform excellently in generating singing voices. However, their performance tends to deteriorate when encountering out-of-distribution data. To handle this, StyleSinger [zhang2023stylesinger] introduces a Residual Style Adaptor and an Uncertainty Modeling Layer Normalization to handle this.
Dataset Language #Singers Hours
NUS-48E [duan2013nus] English 12 1.91
NHSS [sharma2021nhss] English 10 7
JVS-MuSiC [tamaru2020jvs] Japanese 100 2.28
Tohoku Kiritan [ogawa2021tohoku] Japanese 1 1
PopCS [liu2022diffsinger] Chinese 1 5.89
OpenSinger [huang2021multi] Chinese 66 50
Opencpop [wang2022opencpop] Chinese 1 5.25
M4Singer [zhang2022m4singer] Chinese 20 29.77
@@
AMASS [mahmood2019amass] Text
@@
HumanML3D [guo2022generating] Text
@@
Motion-X [lin2024motion] Text
@@
Speech2Gesture [ginosar2019learning] Audio
@@
Talkshow [yi2023generating] Audio
@@
Table 1: Comparison of Audio and Motion Datasets. (a) compares our Rap-Vocal Subset with existing singing vocal datasets. (b) compares our Rap-Motion Subset with existing motion datasets. : “Whole” means whole-body motion including face, body, and hand.
@@
2.2 Text to Motion Generation
Text-to-motion Dataset. Current text-motion datasets, such as KIT [plappert2016kit], AMASS [mahmood2019amass], and HumanML3D [guo2022generating], are constrained by limited data coverage, typically spanning only tens of hours and lacking whole-body motion representation. To overcome these shortcomings, Motion-X [lin2024motion] was developed, providing an extensive dataset that encompasses whole-body motions.
Text-to-motion Models. Text-to-motion models [petrovich2022temos, zhang2024motiondiffuse, guo2022tm2t, guo2022generating, lu2023humantomato, chen2023executing] have gained popularity for their user convenience. Recent advancements have focused on diffusion models [zhang2024motiondiffuse, kim2023flame, tevet2022human], which, unlike deterministic generation models [ahuja2019language2pose, ghosh2021synthesis], enable finer-grained and diverse output generation. Chen et al. [chen2023executing] proposed a motion latent-based diffusion model to enhance generative quality and reduce computational costs. However, the misalignment between natural language and human motions presents challenges. Lu et al. [lu2023humantomato] introduced the Text-aligned Whole-body Motion generation framework to address these issues by generating high-quality, diverse, and coherent facial expressions, hand gestures, and body motions simultaneously.
2.3 Audio to Motion Generation
Audio-to-motion Dataset. Co-Speech Datasets can be classified into 1: pseudo-labeled (PGT) and 2: motion-captured (mocap). PGT datasets [cao2017realtime, habibie2021learning, ginosar2019learning, yi2023generating] utilize disconnected 2D or 3D key points to represent the body, allowing extracting at a lower cost but with limited accuracy. On the other hand, mocap datasets provide annotations for limited body parts, with some focusing solely on the head [fanelli20103, cudeiro2019capture, wuu2022multiface] or the body [takeuchi2017creating, ferstl2018investigating]. Differing from them, BEATX [liu2023emage] contains mesh data of both head and body. However, these datasets typically focus on speech to motion generation. Our RapVerse, in contrast, contains paired text-vocal-motion data, enabling simultaneous motion and vocal generation.
Audio-to-motion Models. Audio-to-motion models [taylor2017deep, shen2023difftalk, zhang2023sadtalker, liu2023emage, yi2023generating, alexanderson2023listen, bian2024adding, chen2024diffsheg] aim to produce human motion from audio innputs. Recognizing the intricate relationship between audio and human face, TalkSHOW [yi2023generating] separately generates face parts and other body parts. However, this approach exhibits limitations, such as the absence of generating lower body motion. Motioncraft [bian2024adding] proposes a DiT structure for motion generation. EMAGE [liu2023emage] utilizes masked gestures together with vocals for a generation. Our model, however, goes a step further by being the first to generate paired audio-motion data directly from text.
@@ 3.1 Rap-Vocal Subset
Vocal and Background Music Separation. Since the crawled songs are mixed with rapping vocals and background music, and we aim to synthesize singing vocals from separated clean data, we utilize Spleeter [hennequin2020spleeter], the state-of-the-art open-source vocal-music source separation tool to separate and extract rap vocal voices and accompanying background music from the collected songs. Following [ren2020deepsinger], we normalize the loudness of the vocal voices to a fixed loudness level.
@@ 3.2 Rap-Motion Subset
The Rap-Motion subset contains 26.8 hours of rap performance videos with 3D holistic body mesh annotations in SMPL-X parameters [pavlakos2019expressive], synchronous singing vocals, and corresponding lyrics. We introduce the collection pipeline of this subset as follows.
@@
Audio Data Processing. Similar to the Rap-Vocal subset, we leverage Spleeter [spleeter2020] to isolate singing vocals from accompanying music. Given that YouTube videos typically lack paired lyrics, we employ an ASR model, Whisper [radford2023robust], to accurately transcribe vocals into corresponding text.
Video Data Processing. To ensure the collection of high-quality video clips for motion annotation, we implemented a semi-automatic process to filter out undesirable content, such as advertisements, transition frames, changes in shots, and flashing lights. Initially, we applied YOLO [redmon2018yolov3] for human detection to discard frames where no humans were detected. Subsequently, we utilized RAFT [teed2020raft] to assess the motion magnitude, employing a threshold to eliminate frames affected by camera instability. We then perform meticulous manual curation on the extracted clips, retaining only those of the highest quality. Finally, we follow the pipeline of the optimized-based method Motion-X [lin2024motion] to extract 3D whole-body meshes from monocular videos. Specifically, we adopt the SMPL-X [pavlakos2019expressive] for motion representations, given a T-frame video clip, the corresponding pose states $\mathcal{M}$ are represented as:
$\mathcal{M}=\{\mathcal{M}_{f},\mathcal{M}_{b},\mathcal{M}_{h},\zeta,\epsilon\}$ (1)